OpenStudy (anonymous):

can someone help solve this by substitution? Whenever I do it the x's or y's always cancel eachother out. 2x-y=5 -2x+y=7 show by steps plz i dont just want the answer.

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

thats what i simply put?

OpenStudy (anonymous):

that is because they are parallel

OpenStudy (anonymous):

Well,\[-2x+y=-(2x-y)\] so you are trying to show that 5 = -7

OpenStudy (anonymous):

1: y=-5+2x 2:-2x+y=7 subs. 1 into 2 -2x+(-5+2x)=7 -2x-5+2x=7 -5=7